Please help me solve this percentage question: To arrive on time a bus needs 18 hours to complete a normal trip. If the bus must arrive in 10 hours, by what percent must the speed of the bus be increased ? Kindly show the steps.
Answer by ankor@dixie-net.com(22740)

To arrive on time a bus needs 18 hours to complete a normal trip.
If the bus must arrive in 10 hours, by what percent must the speed of the bus be increased ?
:
Let s = the normal speed of the bus
let x = per cent increase in decimal form (for example 1.20 is a 20% increase)
:
distance = 18s (time*speed)
the same distance can be written
dist = 10xs
therefore
10xs = 18s
divide both sides by 10s
x = 1.8, the normal speed has to multiplied by 1.8
That's a 80% increase in speed to make the trip in 10 hrs
:
:
You can confirm this say, normal speed is 50 mph
then dist:
18*50 = 900 mi
increase the speed by 80%: 1.8 * 50 = 90 mph
10*90 = 900 mi